### v1

@@ A.6.1 Strict detection unit association
The first and most obvious refinement is to prevent some detection units from being associated with targets that are too different from their size prior. We do this by defining a ”strict association range” $S_{ar}$ that represents the number of detection units that are available to a target object of a given size. For example, with $S_{ar}=3$, each target can only be associated with the three detection units with the closest theoretical size priors. In practice, we do this by masking elements in the matching score matrix before starting the association process (A.3 frame in Fig. 4). The association is still based on the best current prediction but only using a subset of detection units. Interestingly, we can see the classical YOLO-V2 association process as a special case where $S_{ar}=1$.
For this selection to work, we must define which detection units best suit a given target. For this reason, we implemented three different theoretical proximity measurements. The first one is identical to YOLO-V2 and uses a zero-centered IoU comparison. The second one uses the Euclidean distance between the target and the prior of each detection unit in the 2D box-size space. The third one searches for the minimum absolute value for the target size by comparing what would be $\hat{o}_{w}$ and $\hat{o}_{h}$ for each detection, which is equivalent to measuring a distance in the exponentiated and prior-scaled size space. Our preference goes to the last definition as it is more directly related to the loss expression (Eq. 7). The classical zero-centered IoU was mainly kept so our method can emulate a classical YOLO association.
In the specific context of multiple small objects in the same grid cell, we might need to define multiple detection units with an identical size prior, as we did in Sect. 3.7.2. If that is the case, we consider that all detection units with identical priors are made available together and only count for a single $S_{ar}$ increment. For example, with a list of size-priors $p_{w}=p_{h}=\left[6,6,6,6,12,24\right]$ and $S_{ar}=2$, a square target of size $\hat{w}=\hat{h}=7$ could be associated to all the detection units except the one with the largest size prior. All the detection units with $p_{w}=p_{h}=6$ are considered to represent the same scale and account for the first $S_{ar}$ increment, and the second one is the single detection unit with $p_{w}=p_{h}=12$. With this definition, we can discriminate regarding size scales, with some of them being represented by more possible detection units than others. Ultimately, we expect detection units with identically sized priors to specialize based on criteria other than size or shape. By construction, we define that a value of $S_{ar}=0$ implies that this refinement is turned off and that all detection units are available.
@@ A.6.2 Startup and forced random association
At the beginning of the training, when all the weights are still close to their initialization values, the network can be very sensitive to gradient explosion. In addition, the network can be prone to solution collapse, with only a few detection units actively trying to represent all the targets. Due to our prediction-aware association, it is more difficult to reintegrate the unconstrained detection unit as they are less likely to represent a good match than the already constrained ones. One solution to mitigate this issue is to have a low starting learning rate that progressively increases (Sect. 3.8). However, it works best when combined with a second solution, which is to deactivate the prediction-aware association for a given number of iterations $N_{rand}$ and perform a random target-prediction matching instead (B.2.1 frame in Fig. 4). With this method, each detection unit sees a variety of examples representative of the full training dataset distribution, which constrains them to a range of plausible values while preserving enough diversity to prevent solution collapse. These solutions are already part of some versions of the classical YOLO implementation.
For both the default YOLO and our adapted method, having our detection units specialize in some categories of objects is an efficient way to distribute the network expressivity. However, having mostly independent detection units might result in discontinuities in the predicted values for some object properties. This effect can be striking for additional parameters predicted with the formalism presented in Sect. 2.4. For example, when predicting a source flux, there is an overall scaling with the source size. Consequently, if the detection units are mostly distributed over the object size, they can only learn the corresponding subset of the flux relation. Having access to only a sub-part of the parameter space to be predicted can lead to poor constrain of the underlying function, especially on the edges of the sub-part interval.
Our method can ensure some representation overlap between our detection unit directly from the permissive association process when $S_{ar}>1$. However, this approach only allows representation overlap and does not guarantee it. As for random startup, forcing random associations would allow each detection unit to access parts of the underlying function outside its usual specific parameter sub-space. From this, we added the possibility of a random association at a user-defined rate $\alpha_{rand}$ (also B.2.1 frame in Fig. 4), which ensures better continuity of predicted parameters across all detection units. A side effect of this random association is that it acts as an additional training regularization and helps prevent solution collapse. A very low value for $\alpha_{rand}$ around one or two percent is usually enough to obtain a strong regularization effect. While this solution mitigates the apparition of discontinuities, it can still be observed for some of the parameters represented in Fig. 19 from Sect. 4.4, especially on $Bmaj$ and $Bmin$.
@@ A.6.3 Forced specific detection unit association
For the same reasons we added the strict detection unit association (Appendix A.6.1), there are specific cases where stronger constraints can be beneficial. The first forced association is for small objects that are difficult to identify or position because their features are often badly resolved. Also, for small objects, higher precision is required on the positioning to achieve a given prediction-target overlap, which also means that the chances of a random match are low. For these reasons, it is often useful to force the association of very small objects with the detection unit with the smallest size prior in contexts where $S_{ar}>1$. Interestingly, this idea is also present in the YOLO-V1 but was abandoned after switching to the fully convolutional formalism in YOLO-V2. In our method, this forced association is made by comparing the target surface to a limit value that is defined as $\alpha_{small}{\times}s_{w}{\times}s_{h}$, with $s_{w}$ and $s_{h}$ the width and height of the smallest size-prior, and $\alpha_{s}$ a user-defined scaling factor (B.2.2 frame in Fig. 4). If multiple detection units have an identical size-prior that is the smallest, they are all considered viable options. The new target-prediction pair is selected by searching for the best current prediction in this subset of detection units. This is equivalent to having a special value of $S_{ar}=1$ for very small targets, regardless of the global value of $S_{ar}$.
The second forced association searches for the best theoretical detection unit. In our default association process, ordering the association following the matching score matrix helps to associate predictions with the target object they are actively trying to represent. On the contrary, this process is not helpful when a target object is not currently represented by any of the remaining predictions placed randomly in the background. For illustration, in Fig. 26, target A does not have any good association, and the choice to associate it with prediction 2 rather than prediction 1 is solely based on distance. While it can be a good criterion to select the current best prediction when it is actually trying to represent the object, it becomes less informative with predictions that are so far away that they do not represent anything and are likely random. In such a case, it might be better to associate the target with the available detection unit that should theoretically be the best suited to represent it. In this regard, the classical YOLO association, which is based solely on this principle, can be considered a good ”default” association when no assumption can be made regarding each prediction’s relevance other than its prior size. In our method, this forced association with the best theoretical detection unit is adopted when ${\rm fIoU}\leq L^{{\rm fIoU}}_{low}$ with ${\rm fIoU}$ the score of the current target-prediction pair and $L^{{\rm fIoU}}_{low}$ a user-defined threshold. When this refinement is triggered, it will search for the best theoretical prior as described in Appendix A.6.1 accounting for multiple identical priors and excluding already locked boxes (B.2.3 frame in Fig. 4). In addition, this refinement can also serve as a regularization by defining a user-defined rate $\alpha_{best}$ at which this refinement is applied regardless of the ${\rm fIoU}$ value for the current target-prediction pair. This specifically helps prevent solution collapse in cases where $S_{ar}$ is large by forcing all detection units to be used at some point. We found that $\alpha_{best}=0.05$ works well as a regularization for datasets based on everyday life images.
@@
This last refinement can also be used in another way. By setting it to a high value up to $\alpha_{best}=0.95$ in association with a large $S_{ar}$, we are emulating a case where $S_{ar}=1$ but allowing another detection unit that is currently better at detecting a given object to get the association at a low $1-\alpha_{best}$ rate. This type of setup is especially useful for contexts with a strongly imbalanced scale distribution in the training sample, and it is also better at improving the continuity of the predicted parameters across the available detection units. We stress that this approach is very different from a setup with $S_{ar}=1$ and a low $\alpha_{rand}$ rate for which the quality of the current prediction would never matter.
@@ A.6.4 Difficult flagging and quality check
All target objects are not equally difficult to detect, be it due to specific class features, apparent size, strong context dependency, partial obscuring from another object, or proximity to the edges of an image. Moreover, the detection difficulty is not a discrete property. From this, deciding if an object should be included as a target in the training sample regarding a specific application and set of objectives is complex. However, there are cases where both removing or keeping the object can degrade the network performance. For an object only partially visible on the edge of an image, the network might not always be capable of detecting it depending on the apparent features and the current network state. As exposed in Sect. 3.4, if we choose to remove the object while the network is confident in predicting its presence, it will be tasked to lower the objectness score of all objects with similar detectable features. In the opposite case, keeping an object while the network cannot identify any relevant feature will add a lot of noise to the training. This scenario is actually common, especially when performing data augmentation. A classical systematic approach for handling objects close to an edge is to keep them as targets if at least 50% of their expected size is visible in the input data. With this approach, the capability of the network to detect the object can strongly depend on which half is visible. For a target labeled as a cat, having its face visible with the associated features would be more valuable than having just its backside, which could be confused with the likely many similar classes in the training sample.
@@
Several reasons could make an object difficult but not impossible to detect. For all these strongly application-dependent cases, we designed a ”difficult flag” that can be attributed to any target object and for which we apply a ”positive reinforcement only” update. Any target with this flag activated is considered a match if the network is confident enough in its current prediction. This is controlled by two user-defined thresholds based on the predicted box quality ${\rm fIoU}\geq L^{{\rm fIoU}}_{diff}$ and current confidence $O\geq L^{Obj}_{diff}$ of the network (B.2.4 frame in Fig. 4). If one of these limits is not fulfilled, we consider there is no match, but the target still counts for defining the GBNB masks. It means that a good match with low confidence is not penalized, allowing its confidence to increase with further training from similar targets. These limits are tested after all the association refinements, so if the quality condition is not fulfilled, the associated detection unit is considered free for subsequent associations.
@@
Difficult objects are obviously difficult to detect but can also be difficult to classify or characterize. For example, predicting the flux of an astronomical source cut on the edge of the image requires the network to learn symmetrical assumptions that are only required and constrained for this specific context. It is sometimes better to rely on a confident target to properly define the interesting features and to only try to detect or characterize difficult objects based on resemblance. To encapsulate these subtleties, our difficult flag can take different values controlling which parameters get updated when the object passes the quality check: ”0” when the object is not difficult; ”1” when all the parameters should be updated; ”2” when the box geometry, objectness, and probability should be updated, but not the class nor the extra parameters; and ”3” when only the objectness and probability should be updated. All these flagging only affect training and are not reflected in the prediction outside the general objectness score.
@@ A.7 YOLO-CIANNA cascading loss
In sect. 2.7, we introduced the concept of cascading loss, which refers to the fact that the loss varies during training as a function of the current state of the network prediction. This is a direct consequence of having a prediction-aware association in addition to various quality checks on the loss sub-parts and multiple association refinements. In practice, at the beginning of the training, the loss should be dominated by position and size updates toward the most obvious target objects. As the network improves for these objects, their respective loss term will lower, making the contribution from less obvious objects more dominant. Simultaneously, the obvious objects will become sufficiently well detected for other aspects like the objectness, the class, or the additional parameters to be included. From the loss perspective, it results in a competition between better characterizing the obvious objects and starting to detect less obvious ones.
In summary, the better the network gets, the more complete and complex the loss becomes. The objective is to order the network expressivity toward the most important aspects at a given point in the training, not bothering to adjust currently irrelevant properties. Looking at the evolution of the validation loss during training, we usually observe that the loss for the parameters with a quality threshold starts low, then increases before lowering again and reaching a first plateau. This behavior tends to repeat several times during the training with a lower increase and subsequent plateau with a smaller value each time, hence the ”cascading” name. The same behavior can also appear on objectness and probability independently of the fitting limits due to possible regime change in the object distribution over the detection units or simply due to the ”difficult” flagged objects being progressively integrated into the loss.
However, it becomes difficult to identify overtraining with a loss that does not monotonously decrease. To counter this, our implementation contains a parameter to select if the monitored loss on the validation datasets should include all the refinements and limits (”complete” loss) or not (”natural” loss). In the second case, only the base association process is maintained, all difficult objects are ignored, and all the loss parameters are computed for every association regardless of their quality. This usually results in a more classical loss behavior suitable for monitoring overtraining while keeping the cascading loss for the training dataset only. In Fig. 13, we use the ”natural” loss, so the cascading effect is not visible in the monitored validation loss.
@@ A.8 Classical detection dataset benchmark
This section presents a quick summary of the results we obtained on classical object detection datasets using our YOLO-CIANNA method. This aims to validate that our implementation is at least on par with the classical YOLO-V2 performances for everyday images and that our modifications and refinements in the context of astronomical images have not impaired the capabilities of our method for other applications. Since our objective is to evaluate the effect of our low-level modifications, we tried to reproduce the darknet-19 architecture from YOLO-V2 as closely as possible. The later darknet-53 from YOLO-V3 is currently out of reach with CIANNA. We then present classical benchmark results of our slightly modified architecture on the ImageNet, Pascal VOC, and COCO datasets. We only present the most critical aspects of these benchmarks because analyzing them is not the scope of this paper. Still, all the corresponding scripts with all the parameters are provided on the CIANNA git repository and archived with the V1.0 release.
A.8.1 YOLO-CIANNA structural limitations
Our CIANNA framework can already build and train various CNN architectures. However, we still miss some currently widespread structural elements and optimization techniques. Indeed, the overall complexity and diversity of methods in the global ML research field pushed us to implement functionalities and capabilities regarding their usefulness for our specific subset of astronomical applications. We describe here two features that are missing to reproduce the YOLO-V2 and V3 architectures exactly, namely the batch normalization and the skip connections.
While network expressivity, and therefore predictive performance, is supposed to increase with the number of layers, there is a point where it becomes counterproductive to add more layers due to vanishing gradient issues during the training process. This problem is long known (Bengio et al. 1994), and many proposed improvements of deep neural network structures over the last decades were directly aimed at mitigating this issue (e.g., the rectified linear unit activation, Xu et al. 2015). One of the widely adopted mitigation techniques is based on inter-layer normalization. The most common is the batch normalization (BN, Ioffe & Szegedy 2015) that renormalizes the activation map elements of a given filter over the full current batch to have a common mean and variance that is learned during training. It homogenizes the training between layers, strongly mitigating the vanishing gradient issue, and it also regularizes the network over the batch content during training. While the first effect would be beneficial in our case to allow for deeper architecture, the second one pushes the network to ignore the absolute value of the activation map content and rather works only on relative contrast. We expected this to be at least partly detrimental for astronomical images with a large dynamic or when trying to predict pixel-value-dependent parameters like the source flux.
For this reason, the implementation of BN was postponed, and we rater focused on implementing a known alternative technique called group-normalization (GN, Wu & He 2018). It works by normalizing over the activation map elements of a given group of filters but independently for each image in a batch. It preserves the vanishing gradient mitigation principle while removing the regularization aspect. It was demonstrated that replacing all BN operations with GN usually leads to similar performances on many computer vision tasks (Wu & He 2018). Contrary to BN, it comes with an independent hyper-parameter, the group size $g_{s}^{l}$, that defines the number of filters $s$ to be normalized together at a giver layer $l$. Interestingly, implementing GN provides access to two other classical layer normalization methods that are specific cases of GN, the layer normalization with $g_{s}^{l}=N_{f}$ (Lei Ba et al. 2016) and the instance normalization with $g_{s}^{l}=1$ (Ulyanov et al. 2017).
Another typical way to construct deeper networks is to build skip-connections or residual connections (He et al. 2015). These two approaches counter the vanishing gradient problem by adding shortcuts between the layers. In both cases, the shortcut usually spans over a few layers considered as a block, and very deep architectures can be built by stacking several blocks. The network will be first trained as a shallower network to propagate the error up to the higher layers and then progressively get more layers involved to refine its solution with a higher expressivity. Currently, this process is missing from CIANNA, which limits the network depth for some applications. Still, we have not yet encountered an astronomical application for which this is the clear main limiting factor. The absence of skip-connections explains why we cannot reproduce large architectures like darknet-53 from YOLO-V3. In this regard, our emulation of the darknet-19 from YOLO-V2 is also a simplification as we miss the one skip-connect near the end of the network, which can have a significant impact on performances for small objects (Appendix A.8.5).

### v2

@@ A.6.1 Strict detection unit association
The first refinement is to prevent some detection units from being associated with targets that are too different from their size prior. We do this by defining a “strict association range” $S_{ar}$ that represents the number of detection units available to a target object of a given size. For example, with $S_{ar}=3$, each target can only be associated with the three detection units with the closest theoretical size priors. In practice, we do this by masking elements in the matching score matrix before starting the association process (A.3 frame in Fig. 5). The association is still based on the best current prediction but uses only a subset of detection units depending on each target.
For this selection to work, we must define which detection units best suit a given target. We implemented three different theoretical proximity measurements. The first uses a zero-centered IoU comparison like the classical YOLO. The second one uses the Euclidean distance between the target and the prior of each detection unit in the 2D box-size space. The third one searches for the minimum absolute value for the target size by comparing what would be $\hat{o}_{w}$ and $\hat{o}_{h}$ for each detection, which is equivalent to measuring a distance in the exponentiated and prior-scaled size space. Our preference goes to the third approach as it is more directly related to the loss expression (Eq. 7). The classical zero-centered IoU was mainly kept so our method can emulate a classical YOLO association by setting $S_{ar}=1$.
In case identical priors are defined, they are all made available and only count for a single $S_{ar}$ increment. For example, with a list of size-priors $p_{w}=p_{h}=\left[6,6,6,6,12,24\right]$ and $S_{ar}=2$, a square target of size $\hat{w}=\hat{h}=7$ could be associated with all the detection units except the one with the size prior of 24. With this definition, we can distribute the target in different size categories while allowing the detection of multiple objects in the same grid cell for specific size categories. Ultimately, we expect detection units with identical size priors to specialize based on other criteria. By construction, we define that a value of $S_{ar}=0$ turns off this refinement with all detection units available.
@@ A.6.2 Startup and forced random association
At the beginning of the training, when all the weights are still close to their initialization values, the network can be exposed to gradient divergence. One solution to mitigate this issue is to have a low starting learning rate that progressively increases (Sect. 3.8). However, it works best when combined with a second solution that is to deactivate the prediction-aware association for a given number of iterations $N_{\rm rand}$ in favor of a random target-prediction matching (B.2.1 frame in Fig. 5). This refinement is notably found in specific versions of the classical YOLO. It allows each detection unit to see a variety of examples representative of the full training dataset distribution, constraining them to a permissive range of plausible values while preserving diversity.
One other problem we can face with a prediction-aware association is solution collapse, with only a subset of detection units trying to represent all the targets. It reduces the capability of the detector to detect multiple objects, resulting in a suboptimum distribution of the network expressivity. While this issue mostly appears at the early stage of training, it is likely to persist indefinitely as unconstrained detection units are less likely to be a good match. To prevent solution collapse, we need to push detection units to specialize in mostly independent categories of objects. However, this might result in discontinuities in the predicted values for some object properties. For example, as detection units distribute over size categories, there will likely be intermediate-size regimes for which none of the available detection is accurate as it is on the edge of the representation space for all of them. Overall, some degree of specialization of the detection units is required, but it must remain permissive enough so each detection unit can sample a wider parameter space.
Using a strict size association refinement with a permissive range $S_{ar}>1$ already provides some leverage over the permissivity versus specialization balance. However, this approach only allows representation overlap and does not guarantee it. As for random startup, forcing random associations at a user-defined rate $\alpha_{\rm rand}$ can force each detection unit to occasionally access parts of the underlying function outside its usual specialized parameter subspace, ensuring better continuity (also B.2.1 frame in Fig. 5). A very low $\alpha_{\rm rand}$ value around one or two percent is usually enough to prevent mode collapse and to obtain a strong additional regularization effect. However, regarding parameter discontinuity, random association only mitigates the effect that remains measurable in many application cases.
@@ A.6.3 Forced specific detection unit association
There are some specific cases where more constraints in the association can be beneficial. The first forced association is for small objects, which are often difficult to identify or position due to poorly resolved features. In addition, getting small boxes to overlap requires a higher positioning accuracy, and the chances of random matches in the association process are reduced compared to larger boxes. To overcome these issues, we force the association to the smallest prior if the target surface is smaller than a limit value defined as $\alpha_{\rm small}{\times}s_{w}{\times}s_{h}$, with $s_{w}$ and $s_{h}$ the width and height of the smallest size-prior, and $\alpha_{s}$ a user-defined scaling factor (B.2.2 frame in Fig. 5). If multiple detection units have an identical size-prior that is the smallest, they are all considered viable options. The new target-prediction pair is then selected by searching for the best current prediction in this subset of detection units.
The second forced association is for targets that no detection units actively try to identify. In such a context, associating the closest prediction results in selecting a random detection unit most of the time. For example, in Fig. 26, target A has no good association, and the two remaining available predictions fall far away in the background. In such a case, instead of using the distance, it might be better to associate the target with the available detection unit that should theoretically be the best suited for it. In our method, this forced association with the best theoretical detection unit is adopted when ${\rm fIoU}\leq L^{{\rm fIoU}}_{\rm low}$ with ${\rm fIoU}$ the score of the current target-prediction pair and $L^{{\rm fIoU}}_{\rm low}$ a user-defined threshold. The best theoretical detection unit is defined as described in Appendix A.6.1 accounting for multiple identical priors and excluding already locked boxes (B.2.3 frame in Fig. 5).
@@
Finally, we can also use this principle as a regularization approach by defining an $\alpha_{\rm best}$ rate at which the best theoretical detection unit is used regardless of the ${\rm fIoU}$ value for the current target-prediction pair. This is another way to mitigate solution collapse in cases where $S_{ar}$ is high. We found that $\alpha_{\rm best}=0.05$ works well as a regularization for datasets based on everyday life images. This last approach can also be used in an alternative way. By setting it to a high value, up to $\alpha_{\rm best}=0.95$, in association with a high $S_{ar}$, we are emulating a case where $S_{ar}=1$ but allowing another detection unit for which the current prediction is better to get the association at a low $1-\alpha_{\rm best}$ rate. This type of setup is especially relevant for contexts with a strongly imbalanced scale distribution in the training sample, and it is also better at improving the continuity of the predicted parameters across the available detection units. We stress that this approach is different from a setup with $S_{ar}=1$ and a low $\alpha_{\rm rand}$ rate for which the quality of the current prediction would never matter.
@@ A.6.4 Difficult flagging and quality check
All target objects are not equally difficult to detect, be it due to specific class features, apparent size, strong context dependency, partial obscuring from another object, or proximity to the edges of an image. Moreover, the detection difficulty is not a discrete property. From this, deciding if an object should be included as a target in the training sample regarding a specific application and set of objectives is complex. There are even cases where both removing or keeping the object can result in poor performance. As stated in Sect. 3.4, removing a target that the network would have confidently detected will lower the objectness score of all objects with similar detectable features. In the opposite case, keeping a target that is impossible to detect for the network will add noise to global objectness score training. Several reasons could make an object difficult but not impossible to detect. A typical example arises when doing image augmentation, where objects can be cut on the edge of a zoomed-in image. In such a case, the detectability of the object can depend strongly on what part of the object remains, making criteria based on box size fraction inefficient.
@@
To help solve this issue, we designed a “difficult flag” that can be attributed to any target object and for which we apply a “positive reinforcement only” update. Flagged targets can only be matched with predictions for which the network is confident in its prediction. This is controlled by two user-defined thresholds based on the predicted box quality ${\rm fIoU}\geq L^{{\rm fIoU}}_{\rm diff}$ and current confidence $O\geq L^{\rm obj}_{\rm diff}$ of the network (B.2.4 frame in Fig. 5). If one of these limits is not fulfilled, we consider there is no match, but the target still counts for defining the GBNB masks. It means that a good match with low confidence is not penalized, allowing its confidence to increase with further training from similar targets. These limits are tested after all the association refinements, so if the quality condition is not fulfilled, the associated detection unit remains free for subsequent associations.
@@
Objects can not only be difficult to detect but also to classify or characterize. For example, predicting the flux of an astronomical source only partially visible on the edge of an image requires the network to learn symmetrical assumptions. To allow more control, our difficult flag can take different values defining which parameters get updated when the object passes the quality check: “0” when the object is not difficult; “1” when all the parameters should be updated; “2” when the box geometry, objectness, and probability should be updated, but not the class nor the extra parameters; and “3” when only the objectness and probability should be updated. All these flagging only affect training and do not impact the prediction process directly but reflect in the predicted objectness score.
@@ A.7 YOLO-CIANNA cascading loss
In sect. 2.7, we introduced the concept of cascading loss, which refers to the loss changing during training as a function of the current state of the network prediction. This is a direct consequence of having a prediction-aware association in addition to various quality checks on the loss subparts and multiple association refinements. In practice, at the beginning of the training, the loss should be dominated by position and size updates toward the most obvious target objects. As the network improves for these objects, their respective loss term will lower, allowing terms from less obvious objects to be more dominant. Simultaneously, the obvious objects will become sufficiently well detected for other aspects like the objectness, the class, or the additional parameters to be included. From the loss perspective, it results in the search for a balance between better characterizing the most striking objects and starting to detect less obvious ones.
In summary, the better the network gets, the more complete and complex the loss becomes. The objective is to guide the network expressivity toward the most important aspects at a given point in the training, not bothering to adjust currently irrelevant properties. Looking at the evolution of the validation loss during training, we usually observe that the loss for the parameters with a quality threshold starts low, then increases before lowering again and reaching a first plateau. This behavior tends to repeat several times during the training with a lower increase and subsequent plateau with a smaller value each time, hence the “cascading” name. The same behavior can also appear on objectness and probability independently of the fitting limits due to possible regime change in the object distribution over the detection units or simply due to the “difficult-flagged” objects being progressively integrated into the loss.
Identifying overtraining with a dynamic loss is challenging. To counter this, our implementation can monitor the validation loss with (“complete” loss) or without (“natural” loss) the refinements and limits. In the second case, only the base association process is maintained when computing the error on the validation set, and all the loss parameters are computed for every association regardless of their quality. This usually results in a more classical loss behavior suitable for monitoring overtraining while keeping the cascading loss for the training dataset only. In Fig. 13, we use the ”natural” loss, so the cascading effect is not visible in the monitored validation loss.
@@ A.8 Classical detection dataset benchmark
This section presents a quick summary of the results we obtained on classical object detection datasets using our YOLO-CIANNA method. This aims to validate that our implementation is at least on par with the classical YOLO-V2 performances and that design choices motivated by astronomical image properties have not impaired the capabilities of the method for other applications. We tried to reproduce the darknet-19 backbone from YOLO-V2 as closely as possible, considering the current limitation of the CIANNA framework. The later darknet-53 backbone from YOLO-V3 is currently out of reach with CIANNA. We then present classical benchmark results on the ImageNet, Pascal VOC, and COCO datasets. We only present the most critical aspects of these benchmarks, as analyzing them is not the scope of this paper. Still, all the corresponding scripts with all the parameters are provided on the CIANNA git repository and archived with the V-1.0 release (10.5281/zenodo.12806324, Cornu 2024b). The different trained models are also available and archived at 10.5281/zenodo.12801421(Cornu 2024a).
A.8.1 Backbone pretraining with ImageNet-2012
Following the description in Redmon & Farhadi (2017), we build a network backbone as similar as possible to the darknet-19 architecture, which comprises 19 convolutional layers and a few pooling layers. The main structural difference is that we use group normalization layers (Wu & He 2018) instead of batch normalization layers (Ioffe & Szegedy 2015), which should not significantly impact our result as long as we compensate for the missing batch scale regularization effect with more image augmentation. We also miss the skip connection near the end of the network, which reduces our ability to detect small objects.
We first train this backbone on the ImageNet-2012 dataset (ILSVRC-2012, Russakovsky et al. 2015) composed of roughly 1.2 million single-label images with 1000 possible classes so it can identify generic low-level features that are likely relevant for various applications. We first trained our network at the typical $224{\times}224$ input resolution up to a loss plateau that achieved a Top-1 accuracy of 70.1% and a Top-5 accuracy of 89.4% on the validation dataset. We then continue to train the network at a $448{\times}448$ input resolution, which is closer to the desired input resolution for our object detectors. With access to finer features, the network achieved a Top-1 accuracy of 74.7% and a Top-5 accuracy of 91.7%. For comparison, the classical darknet-19 architecture reaches a top-1 accuracy of 76.5% and a top-5 accuracy of 93.3%. The remaining difference is likely due to the absence of the regularization effect of batch normalization. When diverging slightly from the original darknet-19 to include more regularization in another form, for example, adding two large layers with 50% dropout, it is possible to bridge the accuracy gap.